Location Score Factor (100 points each) Weight Convenience Parking facilities Display area Shopper traffic Operating costs Neighborhood .15 20 18 .27 10 10 1.00 86 70 86 90 86 90 83 98 94 89 96 84 90 91 86 a. Using the above factor ratings, calculate the composite score for each location. (Do not round intermediate calculations. Round your final answers to 2 decimal places.) Location Composite Score
Answer #1

Answer

Calculation of composite score for each location

Factor Weight A B C
Convenience 0.15 86 77 83
Parking facilities 0.20 70 88 98
Display area 0.18 86 90 94
Shopper traffic 0.27 90 88 89
Operating costs 0.10 86 91 96
Neighborhood 0.10 90 86 84

Formula for composite score =\sumweight * score

There fore composite score is

Factor A B C
Convenience 0.15 * 86 = 12.9 0.15 * 77 = 11.55 0.15 * 83 = 12.45
Parking facilities 0.20 * 70 = 14 0.20 * 88 = 17.6 0.20 * 98 = 19.6
Display area 0.18 * 86 = 15.48 0.18 * 90 = 16.2 0.18 * 94 = 16.92
Shopper traffic 0.27 * 90 = 24.3 0.27 * 88 = 23.76 0.27 * 89 = 24.03
Operating costs 0.10 * 86 = 8.6 0.10 * 91 = 9.1 0.10 * 96 = 9.6
Neighborhood 0.10 * 90 = 9 0.10 * 86 = 8.6 0.10 * 84 = 8.4
Composite score 84.28 86.81 91